OpenStudy (anonymous):

Just answer to a LEQ.. medal and a fan. How can the properties of rational exponents be applied to simplify expressions with radicals or rational exponents?

OpenStudy (johnweldon1993):

well we can write \[\huge \sqrt{2} = 2^{\frac{1}{2}}\] for example...and also \[\huge \sqrt[5]{78}^3 = 78^{\frac{3}{5}}\]

OpenStudy (johnweldon1993):

It can work the opposite way too... \[\huge 6^\frac{10}{2} = \sqrt{6}^{10}\] or \[\huge 2^\frac{1}{7} = \sqrt[7]{2}\]

OpenStudy (anonymous):

how do you put that into words?

OpenStudy (johnweldon1993):

hmm... if you have an expression with a radical..you can simplify by expressing the radical as a rational exponent, likewise, if you have an expression with a radical exponent, you can express the function as a radical

OpenStudy (anonymous):

Thank you so much . I have like 2 more leq's ill just post them lol i dont wanna over work you

OpenStudy (johnweldon1993):

Haha anytime!